An operation manager at an electronics company wants to test their amplifiers. The design engineer claims they have a mean output of 113113 watts with a variance of 100100.

What is the probability that the mean amplifier output would differ from the population mean by less than 2.72.7 watts in a sample of 4343 amplifiers if the claim is true? Round your answer to four decimal places.
Answer #1

P( < 110.3)

= P(( - )/() < (110.3 - )/())

= P(Z < (110.3 - 113)/(10/))

= P(Z < -1.77)

= 0.0384